OpenStudy (anonymous):

Bruce has a bottle that contains 60% of lemon juice and the rest water. The bottle has 1 liter of water.

OpenStudy (anonymous):

Bruce has a bottle that contains 60% of lemon juice and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of lemon juice and water in the bottle. Define the variable used in the equation. (5 points) Part B: How many liters of lemon juice are present in the bottle? Show your work.

OpenStudy (math&ing001):

You're looking for the amout of juice in the bottle and you got: 40% water = 1 liter 60% juice = ? liters 60*1/40= 1.5 liters of juice

OpenStudy (anonymous):

but wheres the veriable?

OpenStudy (math&ing001):

You could use x : representing number of liters of juice So you'll get 60=40x

OpenStudy (anonymous):

so the equation would be 60*1/40=40x?

OpenStudy (math&ing001):

it would be 60*1/40=x

OpenStudy (anonymous):